# SAR-TEXT: A Large-Scale SAR Image-Text Dataset Built with SAR-Narrator and Progressive Transfer Learning

### v1

@@ Abstract
Vision Language Models (VLMs) have achieved remarkable breakthroughs in the field of remote sensing in recent years. Synthetic Aperture Radar (SAR) imagery, with its all-weather capability, is essential in remote sensing, yet the lack of large-scale, high-quality SAR image-text datasets hinders its semantic understanding. In this paper, we construct SAR-Text, a large-scale and high-quality dataset consisting of over 130,000 SAR image–text pairs. To construct the SAR-Text dataset, we design the SAR-Narrator framework, which generates textual descriptions for SAR images through a multi-stage progressive transfer learning strategy. To verify the effectiveness of the SAR-TEXT dataset, we conduct experiments on three typical vision-language tasks: image-text retrieval, image captioning, and visual question answering (VQA). Specifically, we construct three representative models on SAR-TEXT: SAR-RS-CLIP, SAR-RS-CoCa, and SAR-GPT. SAR-RS-CLIP achieves notable improvements in retrieval performance, boosting average recall by 16.43% and 10.54% on the OSdataset-512 and HRSID test sets, respectively. In the captioning task, SAR-RS-CoCa achieves BLEU-4, SPICE, and CIDEr scores exceeding those of the original CoCa model by more than 8$\times$, 4$\times$, and 10$\times$, respectively. In the VQA task, SAR-GPT outperforms baseline and single-stage models on multiple SAR-VQA datasets, demonstrating stronger semantic understanding and reasoning ability, as further confirmed by qualitative results.
It is worth noting that, as a flexible captioning tool, SAR-Narrator can be readily adopted by the community to construct larger-scale SAR image–text datasets.
@@ 1 Introduction
Figure 2: Overview of the SAR-Narrator framework and SAR-TEXT dataset: SAR images are automatically captioned by SAR-Narrator to construct SAR-TEXT, a high-quality dataset containing 130,000+ text-image pairs. Subsequently, two models—SAR-CLIP and SAR-CoCa—are trained using vision-language models (VLMs) and a progressive two-stage fine-tuning strategy to achieve more effective semantic interpretation of SAR images.
@@
However, due to the complex imaging mechanisms of SAR and the substantial differences in visual texture compared to optical imagery, semantic interpretation of SAR data has long presented significant challenges. Traditional approaches often rely on expert-driven manual analysis, which is time-consuming, labor-intensive, and insufficient to meet the increasing demand for efficient SAR image interpretation. Consequently, several studies have attempted to employ machine learning techniques to assist in SAR image interpretation, yet their effectiveness remains limited [46, 1]. With the advancement of deep learning techniques [20, 4], research in remote sensing has made significant progress, particularly through recent developments in vision–language foundation models (VLFMs) such as CLIP [22] and CoCa [35], which have achieved remarkable success in optical remote sensing. However, these advancements remain largely confined to optical imagery, as VLFMs continue to struggle with SAR data due to the substantial modality gap and the scarcity of annotated SAR–text pairs [11]. More notably, the substantial differences between human visual perception and SAR image representations not only increase the complexity of annotation tasks but also necessitate specialized domain knowledge and extensive training for annotators, leading to high annotation costs. Consequently, existing SAR datasets are limited in both scale and diversity. These challenges underscore the importance of constructing large-scale, high-quality SAR image–text annotation datasets and developing efficient algorithms and models specifically designed for SAR imagery.
To address the aforementioned challenges, this paper presents SAR-Narrator, a SAR image caption annotation tool, which is employed to construct a large-scale SAR image–text paired dataset named SAR-TEXT, which contains more than 130,000 SAR images and their corresponding high-quality natural language descriptions, effectively alleviating the problem of scarce annotation data in this field. In addition, a two-stage progressive fine-tuning strategy is proposed to adapt vision–language foundation models, specifically CLIP and CoCa, with the goal of bridging the modality gap between SAR and natural images and enhancing the semantic interpretability of SAR data. Experimental results demonstrate the effectiveness of both the SAR-Narrator tool and the proposed progressive transfer learning approach. Details are shown in Figure 2.
@@
    A progressive fine-tuning strategy is introduced for adapting vision–language foundation models (VLFMs) to SAR imagery. This leads to the development of SAR-RS-CLIP and SAR-RS-CoCa, the first VLFMs specifically trained for SAR image–text understanding. Experimental results demonstrate that both models substantially outperform baseline methods in SAR image–text retrieval and caption generation tasks.
@@ 2.2 VLMs for Remote Sensing
As previously noted, vision–language models (VLMs) have shown significant potential for a wide range of applications in the remote sensing domain. Their use spans multiple tasks, including land cover classification and scene understanding, target detection and object recognition in remote sensing imagery, change detection and disaster assessment, semantic retrieval, and information querying.
@@ 3.1.1 A2C Algorithm
We propose an Annotation-to-Caption (A2C) algorithm designed to transform object detection annotations into descriptive captions. The core steps of the A2C algorithm include parsing the categories, quantities, and spatial locations of objects within SAR imagery, followed by automatically generating structured textual descriptions based on detection boxes (e.g., “The image contains three ships located at the center-right.”). Additionally, a semantic optimization module is introduced, employing a language model to refine sentence structure, ensuring captions are natural and fluent. The A2C algorithm is particularly suitable for target detection datasets, such as MSAR-1.0 [3] and HRSID [30], enriching SAR target recognition tasks with detailed textual annotations. The pseudocode of the A2C algorithm is provided in Algorithm 1.
@@ 3.1.2 SA2C Algorithm
[…]
@@
[…]
@@
[…]
@@
3.1.3 SAR image caption generation
For the SAR-optical image pairing datasets, we hypothesize that although directly annotating SAR images poses significant challenges, generating high-quality captions for their corresponding optical images can serve as an indirect semantic annotation approach. Therefore, our annotation work is conducted based on the fundamental assumption that captions generated for optical images in image pairs can effectively approximate the semantic content of their corresponding SAR images. However, due to substantial differences between SAR and optical imagery in imaging modality, texture characteristics, and semantic representation, directly transferring captions from optical images may introduce semantic bias and descriptive inaccuracies.
To address these challenges, we propose a caption rewriting mechanism leveraging the semantic understanding capabilities of large language models. Initially, the CoCa model is fine-tuned on the HQRS-IT-210K dataset to construct RS-Captioner [7, 8, 21], a caption generator specifically designed for remote sensing images, producing preliminary textual descriptions of optical imagery. Subsequently, a large language model, guided by carefully designed prompts and manually annotated examples through an In-Context Learning (ICL) strategy [5], semantically adjusts and optimizes these initial captions to generate refined natural language descriptions that more accurately reflect the semantic characteristics of SAR imagery.
@@ 3.1.3 SAR image caption generation
To achieve high-quality image caption generation, this paper fine-tunes the CoCa model on the HQRS-IT-210K dataset and trains the RS-CoCa model, named RS-Captioner, which is suitable for remote sensing image scenarios. The model is constructed based on 1.26 million image-text pairs, significantly improving CoCa’s descriptive ability in optical remote sensing images. The high-quality captions generated by RS-Captioner provide a reliable textual foundation for subsequent SAR image caption rewriting based on large language models.
To enhance the modal adaptability of subtitles, this paper further introduces the context learning method of large language models to perform targeted rewriting of optical image subtitles, making them more consistent with the visual semantic features of SAR images. In terms of rewriting strategies, this paper systematically analyzes the differences between SAR images and optical images in imaging mechanisms and visual representation, deriving a set of general principles applicable to the rewriting task. These principles are then converted into guided prompts to support the LLM in achieving precise semantic transfer and language generation under contextual conditions. The specific prompt design for rewriting is as follows
@@ 3.2 SAR-TEXT Dataset Construction
SAR-ship 43819 Ship Based on high-res Gaofen-3 and Sentinel-1 SAR data. Contains 43819 ship chips for target recognition.
OSdataset 2249 - Includes 10,692 optical and SAR image pairs (256×256), with optical from Google Earth and SAR from Gaofen-3 over several cities.
HRSID_JPG 3642 Ship 5604 high-res SAR images and 116951 labeled ships at 0.5m, 1m, and 3m resolution.
QXSLAB_SAROPT 20000 - Contains three ports: Jeddah, Shanghai, Rotterdam. SAR from Gaofen-3, resolution 1m.
SEN12 18094 - Includes 282,384 SAR-optical image pairs covering global regions and four seasons.
optical_sar 8575 - 10,000 SAR-optical image pairs for regional analysis. SAR and optical pairs may be misaligned.
whu-sar-opt 1600 Farmland, City, Village, Water, Forest, Road Covers over 50,000km² with optical (Gaofen-1) and SAR (Gaofen-3) data.
SSDD 1160 Ship SAR ship detection dataset using RadarSat-2, TerraSAR-X, and Sentinel-1. Resolution: 1m–15m.
@@ 3.3 Progressive Transfer Learning Strategy
For training the SAR-RS-CLIP model, the CLIP backbone is first pre-trained on the HQRS-IT-210K dataset to adapt it to optical remote sensing image–text matching tasks. It is then further fine-tuned on the SAR-TEXT dataset to complete the transfer to SAR image–text matching. The SAR-RS-CoCa model follows the same two-stage training procedure, with the CoCa backbone also pre-trained on HQRS-IT-210K and subsequently fine-tuned on SAR-TEXT. Ablation experiments confirm the effectiveness of this progressive transfer strategy.
@@ 4.3 Experimental Results
4.3.1 SAR-CLIP Cross-Modal Retrieval Experiment
The section demonstrates the impact of different CLIP training strategies on model performance in SAR-text cross-modal retrieval tasks. Prior to formal training, this paper first evaluates the initial performance of the original CLIP model and the previously fine-tuned HQRS-CLIP model [x] on SAR images. The experimental results indicate that both the standard CLIP and HQRS-CLIP models exhibit suboptimal cross-modal retrieval performance on SAR images, with an average recall rate of less than 1% on the test set, thereby validating the limitations of the direct transfer strategy in SAR scenarios.
In the transfer learning experiment, this paper first fine-tunes the CLIP model based on the SAR-TEXT dataset to obtain the SAR-CLIP model. Experimental results show that compared to the original CLIP baseline, SAR-CLIP achieves average recall rate improvements of 10.54% and 16.43% on the HRSID and OSdataset_512 test sets, respectively, significantly validating the effectiveness of the SAR-TEXT dataset in image-text alignment tasks. This result further demonstrates that SAR-TEXT can provide rich and reliable semantic pairing information for SAR images, significantly enhancing cross-modal retrieval performance. Additionally, experiments show that the CLIP model can achieve significant performance improvements in SAR image tasks through simple fine-tuning, confirming the strong transferability of the “pre-training-fine-tuning” paradigm in multi-modal remote sensing tasks.
To further explore more optimal transfer strategies, this paper draws on the empirical analysis in [x] and proposes a progressive transfer learning scheme. This scheme first pre-trains the CLIP model using remote sensing optical image text data to adapt it to the modal features of the remote sensing field. Subsequently, the model is further fine-tuned to adapt to SAR image tasks, achieving effective knowledge transfer from natural images to SAR images. The model obtained from this two-stage training strategy is named SAR-RS-CLIP.
@@ 4.3.1 SAR-CLIP Cross-Modal Retrieval Experiment
As shown in Tables 2 and 3, the model trained through a two-stage process achieved an average recall rate improvement of approximately 1% and 0.2% on the HRSID and OSdataset_512 test sets, respectively. Notably, the performance improvements on the i2t-R@1 and t2i-R@1 metrics of the OSdataset_512 test set were particularly significant, reaching approximately 1.2% and 0.5%, respectively. These results indicate that the progressive transfer learning strategy—pre-training the CLIP model using remote sensing optical image data and then fine-tuning it with SAR-TEXT data—effectively improves the model’s cross-modal retrieval performance on SAR images.
@@ 4.3.2 Visualization of image and text embedding based on UMAP
Figures 7 and 8 present the UMAP visualization results of the CLIP and SAR-RS-CLIP models on the HRSID test set. As shown in the figures, neither model forms a clear clustering structure of image categories on this dataset, and the overall distribution of image embeddings remains quite disordered. This phenomenon is primarily due to the high homogeneity of the HRSID dataset: all images depict ship targets with similar appearance and texture features, which makes it challenging for the models to distinguish between them effectively. Nevertheless, despite the limited overall separability, the SAR-RS-CLIP model demonstrates clearer image distribution boundaries compared to CLIP and achieves better separation among image embeddings. This indicates that its capacity to capture fine-grained differences is enhanced to some extent through SAR-TEXT training.
Figures 9 and 10 display the text-image embedding visualizations of the two models on the OSdataset-512 test set. Compared to HRSID, the image samples in OSdataset-512 exhibit greater diversity, resulting in a more dispersed and structurally rich distribution of embeddings in the two-dimensional space. Notably, the SAR-RS-CLIP model significantly improves the distinguishability between images and text on this dataset, with embedding clusters of different categories showing more distinct boundaries and clearer text-image pairing structures. This demonstrates its enhanced capabilities in cross-modal alignment and semantic representation, which can contribute to improved performance in downstream text-image retrieval tasks.
@@
4.3.3 SAR-RS-CoCa Image Caption Generation Test Experiment
Table 4 and Table 5 present the performance of different CoCa models on the HRSID and OSdataset test sets. First, this paper evaluates the zero-shot capability of the pre-trained CoCa model and the RS-CoCa model trained in the preceding section in the SAR image captioning task. The experimental results indicate that the CoCa model, which has not been trained specifically for SAR images, performs extremely poorly in this task, barely able to generate effective captions. This suggests that without the ability to align SAR image and text semantics, the model struggles to generalize to the SAR image domain.
The experimental results show that the performance of the two generative models fine-tuned with SAR text-image data has significantly improved on the HRSID and OSdataset test sets, with improvements in multiple metrics reaching several times or even tenfold increases, fully validating the effectiveness and value of the SAR-TEXT dataset in generative tasks. Additionally, SAR-RS-CoCa outperformed SAR-CoCa on both test sets, further confirming the adaptability and advantages of the two-stage transfer learning strategy in generative tasks.

### v3

# SAR-TEXT: A Large-Scale SAR Image-Text Dataset Built with SAR-Narrator and A Progressive Learning Strategy for Downstream Tasks

@@ Abstract
Vision Language Models (VLMs) have achieved remarkable breakthroughs in the field of remote sensing in recent years. Synthetic Aperture Radar (SAR) imagery, with its all-weather capability, is essential in remote sensing, yet the lack of large-scale, high-quality SAR image-text datasets hinders its semantic understanding. In this paper, we construct SAR-TEXT, a large-scale and high-quality dataset consisting of over 130,000 SAR image–text pairs. To construct the SAR-TEXT dataset, we design the SAR-Narrator framework, which generates textual descriptions for SAR images through a multi-stage strategy. To verify the effectiveness of the SAR-TEXT dataset, we conduct experiments on three typical vision-language tasks: image-text retrieval, image captioning, and visual question answering (VQA). Specifically, we construct three representative models on SAR-TEXT: SAR-RS-CLIP, SAR-RS-CoCa, and SAR-GPT. SAR-RS-CLIP achieves notable improvements in retrieval performance, boosting average recall by 12.97% and 10.0% on the OSdataset_512 and HRSID test sets, respectively. In the captioning task, SAR-RS-CoCa achieves significant improvements over the original CoCa models in terms of BLEU-4, SPICE, and CIDEr scores. In the VQA task, SAR-GPT outperforms baseline and single-stage models on multiple SAR-VQA datasets, demonstrating stronger semantic understanding and reasoning ability, as further confirmed by qualitative results.
It is worth noting that, as a flexible captioning tool, SAR-Narrator can be readily adopted by the community to construct larger-scale SAR image–text datasets. All code, pretrained models, and the SAR-Text dataset are publicly available at: https://github.com/YiguoHe/SAR-TEXT.
@@ 1 Introduction
Figure 2: Overview of the SAR-Narrator framework and SAR-TEXT dataset: SAR images are automatically captioned by SAR-Narrator to construct SAR-TEXT, a high-quality dataset containing 130,000+ text-image pairs. Subsequently, two models—SAR-RS-CLIP and SAR-CoCa—are trained using vision-language models (VLMs) and a progressive two-stage fine-tuning strategy to achieve more effective semantic interpretation of SAR images.
@@
However, due to the complex imaging mechanisms of SAR and the substantial differences in visual texture compared to optical imagery, semantic interpretation of SAR data has long presented significant challenges. Traditional approaches often rely on expert-driven manual analysis, which is time-consuming, labor-intensive, and insufficient to meet the increasing demand for efficient SAR image interpretation. Consequently, several studies have attempted to employ machine learning techniques to assist in SAR image interpretation, yet their effectiveness remains limited [49, 1]. With the advancement of deep learning techniques [22, 4], research in remote sensing has made significant progress, particularly through recent developments in vision language foundation models (VLFMs) such as CLIP [24] and CoCa [38], which have achieved remarkable success in optical remote sensing. However, these advancements remain largely confined to optical imagery, as VLFMs continue to struggle with SAR data due to the substantial modality gap and the scarcity of annotated SAR–text pairs [12]. More notably, the substantial differences between human visual perception and SAR image representations not only increase the complexity of annotation tasks but also necessitate specialized domain knowledge and extensive training for annotators, leading to high annotation costs. Consequently, existing SAR datasets are limited in both scale and diversity. These challenges underscore the importance of constructing large-scale, high-quality SAR image–text annotation datasets and developing efficient algorithms and models specifically designed for SAR imagery.
To address the aforementioned challenges, this paper presents SAR-Narrator, a SAR image caption annotation tool, which is employed to construct a large-scale SAR image–text paired dataset named SAR-TEXT, which contains more than 130,000 SAR images and their corresponding high-quality natural language descriptions, effectively alleviating the problem of scarce annotation data in this field. In addition, a two-stage progressive fine-tuning strategy is proposed to adapt vision–language foundation models, specifically CLIP and CoCa, with the goal of bridging the modality gap between SAR and natural images and enhancing the semantic interpretability of SAR data. Experimental results demonstrate the effectiveness of both the SAR-Narrator tool and the proposed progressive transfer learning approach. Furthermore, additional exploration of downstream VQA tasks has demonstrated the practical applicability of the proposed SAR-TEXT dataset. Details are shown in Figure 2.
@@
    A progressive fine-tuning strategy is introduced for adapting vision language foundation models (VLFMs) to SAR imagery. This leads to the development of SAR-RS-CLIP and SAR-RS-CoCa, the first VLFMs specifically trained for SAR image–text understanding. Experimental results demonstrate that both models substantially outperform baseline methods in SAR image–text retrieval and caption generation tasks.
@@ 2.2 VLMs for Remote Sensing
As previously noted, vision–language models (VLMs) [27, 21, 7, 23, 54, 53, 52] have shown significant potential for a wide range of applications in the remote sensing domain. Their use spans multiple tasks, including land cover classification and scene understanding, target detection and object recognition in remote sensing imagery, change detection and disaster assessment, semantic retrieval, and information querying.
@@ 3.1.1 A2C Algorithm
We propose an Annotation-to-Caption (A2C) algorithm designed to transform object detection annotations into descriptive captions. The core steps of the A2C algorithm include parsing the categories, quantities, and spatial locations of objects within SAR imagery, followed by automatically generating structured textual descriptions based on detection boxes (e.g., “There are 3 ships in the top-left of this image.”). The A2C algorithm is particularly suitable for target detection datasets, such as MSAR-1.0 [3] and HRSID [33], enriching SAR target recognition tasks with detailed textual annotations. The pseudocode of the A2C algorithm is provided in Algorithm 1.
@@ 3.1.2 SA2C Algorithm
category_counts $\leftarrow$ {category: 0 for category in rgb_to_category.values()}
@@
proportions $\leftarrow$ {category: (category_counts[category] / total_pixels) $\times$ 100
@@
filtered_proportions $\leftarrow$ {category: proportion for (category, proportion) in proportions
@@
3.1.3 Rule-Guided LLM-ICL Rewriting for SAR Captioning
For the SAR-optical image pairing dataset, we consider that while directly annotating SAR images presents significant challenges, generating high-quality descriptive text for the corresponding optical images can serve as an effective indirect semantic annotation method, as the descriptive text generated for the optical images can effectively approximate the semantic content of the corresponding SAR images. However, due to substantial differences between SAR and optical imagery in imaging modality, texture characteristics, and semantic representation, directly transferring captions from optical images may introduce semantic bias and descriptive inaccuracies.
To address these challenges, we propose a caption rewriting mechanism leveraging the semantic understanding capabilities of large language models. Initially, the CoCa model is fine-tuned on the HQRS-IT-210K dataset to construct RS-Captioner [8, 9, 23], a caption generator specifically designed for remote sensing images, producing preliminary textual descriptions of optical imagery. The high-quality captions generated by RS-Captioner provide a reliable textual foundation for subsequent SAR image caption rewriting based on large language models. Subsequently, a large language model, guided by carefully designed prompts and manually annotated examples through an In-Context Learning (ICL) strategy [5], semantically adjusts and optimizes these initial captions to generate refined natural language descriptions that more accurately reflect the semantic characteristics of SAR imagery.
@@ 3.1.3 Rule-Guided LLM-ICL Rewriting for SAR Captioning
To enhance the modal adaptability of subtitles, this paper further introduces the context learning method of large language models to perform targeted rewriting of optical image subtitles, making them more consistent with the visual semantic features of SAR images. In terms of rewriting strategies, this paper systematically analyzes the differences between SAR images and optical images in imaging mechanisms and visual representation, deriving a set of general principles applicable to the rewriting task. These principles are then converted into guided prompts to support the LLM in achieving precise semantic transfer and language generation under contextual conditions. The specific prompt design for rewriting is as follows.
@@ 3.2 SAR-TEXT Dataset Construction
SAR-ship 43819 Ship The dataset is composed of 102 images from China’s Gaofen-3 satellite and 108 images from Sentinel-1, containing 43,819 ship chips, each with a resolution of 256×256 pixels, with both range and azimuth dimensions of 256 pixels.
OSdataset 11245 - The dataset includes 2673 pairs of 512×512 pixel registered images and 10692 pairs of 256×256 pixel images. The SAR-TEXT training dataset uses only the annotated training and validation images. Optical images are from Google Earth, and SAR images are from China’s Gaofen-3 satellite with 1-meter resolution in C-band full polarization.
HRSID_JPG 3642 Ship The images are acquired from China’s Gaofen-3 satellite, containing 5,604 high-resolution SAR images and 16,951 annotated ship instances, with resolutions of 0.5 meters, 1 meter, and 3 meters.
QXSLAB_SAROPT 20000 - The SAR images are from China’s Gaofen-3 satellite, and the optical images are from Google Earth. The dataset includes 20,000 pairs of corresponding image tiles collected from three port cities: Jeddah, Shanghai, and Qingdao.
SEN12 18094 - The original dataset is collected from multiple satellites, covering global regions and seasonal variations, containing 282,384 pairs of corresponding image tiles. After filtering, 18,094 pairs of samples were selected for annotation.
optical_sar 8575 - The original dataset contains 10,000 pairs of SAR and optical images for regional analysis. After processing and selection, 8,575 images were used for annotation.
whu-sar-opt 1600 Farmland, City, Village, Water, Forest, Road The optical images in the dataset are from China’s Gaofen-1 satellite, and the SAR images are from China’s Gaofen-3 satellite. The original data consists of 100 pairs of optical and SAR images, each approximately 5556×3704 pixels, covering an area of about 50,000 square kilometers in Hubei Province. After slicing the images using a sliding window, 1,600 pairs of samples were obtained.
SSDD 1160 Ship The original dataset contains 5,604 cropped SAR images, from which 1,160 images were selected for annotation after filtering.
@@ 3.3 Progressive Transfer Learning Strategy
For training the SAR-RS-CLIP model, the CLIP backbone is first continual pretraining on the HQRS-IT-210K dataset to adapt it to optical remote sensing image–text matching tasks. It is then further fine-tuned on the SAR-TEXT dataset to complete the transfer to SAR image–text matching. The SAR-RS-CoCa model follows the same two-stage training procedure, with the CoCa backbone also continual pretraining on HQRS-IT-210K and subsequently fine-tuned on SAR-TEXT. Ablation experiments confirm the effectiveness of this progressive transfer strategy.
@@ 4.3 Experimental Results
4.3.1 SAR-RS-CLIP Cross-Modal Retrieval Experiment
The section demonstrates the impact of different CLIP training strategies on model performance in SAR-text cross-modal retrieval tasks. Prior to formal training, this paper first evaluates the initial performance of the original CLIP model and the previously fine-tuned HQRS-CLIP model [9] on SAR images. The experimental results indicate that both the standard CLIP and HQRS-CLIP models exhibit suboptimal cross-modal retrieval performance on SAR images, with an average recall rate of less than 1% on the test set, thereby validating the limitations of the direct transfer strategy in SAR scenarios.
In the transfer learning experiment, this paper first fine-tunes the CLIP model based on the SAR-TEXT dataset to obtain the SAR-RS-CLIP model. Experimental results show that compared to the original CLIP baseline, SAR-RS-CLIP achieves average recall rate improvements of 10.0% and 12.97% on the HRSID and OSdataset_512 test sets, respectively, significantly validating the effectiveness of the SAR-TEXT dataset in image-text alignment tasks. This result further demonstrates that SAR-TEXT can provide rich and reliable semantic pairing information for SAR images, significantly enhancing cross-modal retrieval performance. Additionally, experiments show that the CLIP model can achieve significant performance improvements in SAR image tasks through simple fine-tuning, confirming the strong transferability of the “pre-training-fine-tuning” paradigm in multi-modal remote sensing tasks.
To further explore more optimal transfer strategies, this paper draws on the empirical analysis in [12] and proposes a progressive transfer learning scheme. This scheme first pre-trains the CLIP model using remote sensing optical image text data to adapt it to the modal features of the remote sensing field. Subsequently, the model is further fine-tuned to adapt to SAR image tasks, achieving effective knowledge transfer from natural images to SAR images. The model obtained from this two-stage training strategy is named SAR-RS-CLIP.
As shown in Tables 2 and 3, The two-stage trained SAR-RS-CLIP model achieved an average recall rate that was 10.98% and 13.13% higher than the baseline CLIP model on the HRSID and OSdataset_512 test sets, respectively, and approximately 1% and 0.2% higher than the single-stage SAR-RS-CLIP model. Notably, the performance improvements on the i2t-R@1 and t2i-R@1 metrics of the OSdataset_512 test set is particularly significant, reaching approximately 1.2% and 0.5%, respectively. These results indicate that the progressive transfer learning strategy—pre-training the CLIP model using remote sensing optical image data and then fine-tuning it with SAR-TEXT data—effectively improves the model’s cross-modal retrieval performance on SAR images.
@@ 4.3.2 Visualization of image and text embedding based on UMAP
Figures 7 and 8 present the UMAP visualization results of the CLIP and SAR-RS-CLIP models on the HRSID test set. Due to the high homogeneity of the HRSID dataset, all images feature ship targets with similar appearance and texture characteristics, which makes it challenging for the model to effectively distinguish between them. As shown in the figure, the image distribution of the CLIP model is more concentrated, suggesting that the images are challenging for the model to distinguish. After training, the image distribution of the two-stage SAR-RS-CLIP model becomes significantly more dispersed, indicating that the model’s ability to differentiate between images in the HRSID dataset has improved to some extent. Nevertheless, despite the limited overall separability, the SAR-RS-CLIP model demonstrates clearer image distribution boundaries compared to CLIP and achieves better separation among image embeddings. This indicates that its capacity to capture fine-grained differences is enhanced to some extent through SAR-TEXT training.
Figures 9 and 10 display the text-image embedding visualizations of the two models on the OSdataset_512 test set. Compared to HRSID, the image samples in OSdataset_512 exhibit greater diversity, resulting in a more dispersed and structurally rich distribution of embeddings in the two-dimensional space. Notably, the SAR-RS-CLIP model significantly improves the distinguishability between images and text on this dataset, with embedding clusters of different categories showing more distinct boundaries and clearer text-image pairing structures. This demonstrates its enhanced capabilities in cross-modal alignment and semantic representation, which can contribute to improved performance in downstream text-image retrieval tasks.
@@
4.3.3 SAR-RS-CoCa Image Caption Generation Test Experiment
Table 4 and Table 5 present the performance of different CoCa models on the HRSID and OSdataset test sets. First, this paper evaluates the zero-shot capability of the pre-trained CoCa model and the RS-CoCa model trained by He et al. [9] in the task of caption generation for SAR images. The experimental results indicate that the CoCa model, which has not been trained specifically for SAR images, performs extremely poorly in this task, barely able to generate effective captions. This suggests that without the ability to align SAR image and text semantics, the model struggles to generalize to the SAR image domain.
The experimental results show that the performance of the two generative models fine-tuned with SAR text-image data has significantly improved on the HRSID and OSdataset test sets, with performance enhancements in multiple metrics reaching up to several times, fully validating the effectiveness and value of the SAR-TEXT dataset in generative tasks. Additionally, SAR-RS-CoCa outperformed SAR-CoCa on both test sets, further confirming the adaptability and advantages of the two-stage transfer learning strategy in generative tasks.